I am a 5 digit number. my tens digit is the sum of my hundreds digit and thousands digit. my tens digit is twice my ones digit. my hundrends digit is 1 more than zero. my thousands digit is seven times my hundreds digit. my ten thousands digit is the same as my ones dgit. what is the mystery number?

Answer by ankor@dixie-net.com(22740)

I am a 5 digit number.
let it = 10000v + 1000w + 100x + 10y + z
:
Write an equation for each statement:
;
my tens digit is the sum of my hundreds digit and thousands digit.
y = x + w
:
my tens digit is twice my ones digit.
y = 2z
:
my hundreds digit is 1 more than zero.
x = 1
:
my thousands digit is seven times my hundreds digit.
w = 7x
:
my ten thousands digit is the same as my ones digit.
v = z
:
Replace x with 1 in each equation that has x
y = w + 1
we know w = 7x, therefore:
w = 7
we know y = w+x: therefore
y = 7 + 1
y = 8
y = 2z
2z = 8
z = 4
we know v=z therefore
z = 4
and
v = 4
:
we have:
v = 4
w = 7
x = 1
y = 8
z = 4
:
47184 is the number